OpenStudy (anonymous):

\[evaluate \int\limits\limits \frac{dx}{4x \sqrt{x ^{2}-16}}\]

OpenStudy (mimi_x3):

didnt you just post that question?

OpenStudy (anonymous):

I have a question for that.

OpenStudy (anonymous):

what should I let???

OpenStudy (anonymous):

let u = 4x???

OpenStudy (anonymous):

and du = 4dx???

OpenStudy (anonymous):

isn't this some inverse trig function?

OpenStudy (mimi_x3):

It's a trig sub. \(x=4sec\theta\)

OpenStudy (anonymous):

what @Mimi_x3 ???

OpenStudy (mimi_x3):

\(dx=4sec\theta tan\theta\)

OpenStudy (mimi_x3):

What's wrong?

OpenStudy (anonymous):

I can't get it. :(

OpenStudy (anonymous):

oh so it is

OpenStudy (anonymous):

what should i let???

OpenStudy (mimi_x3):

\[\frac{1}{4} \int\limits\frac{4\tan\theta \sec\theta}{4\sec\theta\sqrt{16\sec^{2}\theta-16}} => \frac{1}{4} \int\limits\frac{4\tan\theta \sec\theta}{4\sec\theta\sqrt{16(\sec^{2}\theta-1)}} \] => \[\frac{1}{4} \int\limits\frac{4\tan\theta \sec\theta}{4\sec\theta\sqrt{16\tan^{2}\theta}} \]

OpenStudy (mimi_x3):

\[\frac{1}{4} \int\limits\frac{4\tan\theta \sec\theta}{4\sec\theta4\tan\theta} \] maybe..if i didntt make a mistake

OpenStudy (anonymous):

wow what an amazing amount of cancellation! just get the integral of 1

OpenStudy (mimi_x3):

lol yeah; a very easy integral

OpenStudy (anonymous):

which leads me to believe we could do this without a trig sub, if you recall that \[\frac{d}{dx}\sec^{-1}(x)=\frac{1}{x\sqrt{x^2-1}}\] but maybe not

OpenStudy (mimi_x3):

table of standard integrals?

OpenStudy (anonymous):

@satellite73 was right.

OpenStudy (mimi_x3):

lol sorry i went the long way; i didnt know that sec was in the table of standard integrals

OpenStudy (anonymous):

you don't need to do it in standard integral. but it's okay @Mimi_x3

OpenStudy (anonymous):

Thanks @Mimi_x3 and @satellite73 :)

OpenStudy (anonymous):

no it is not the long way if you happen do remember the derivative of inverse secant then i think you can use \(u=\frac{x}{4}\) and get \[\frac{1}{u\sqrt{16u^2-16}}\] or something like that probably constants are wrong